OpenStudy (anonymous):

USE THE GIVEN BASE TO WRITE THE NUMBER AS AN EXPONENTIAL EXPRESSION 9/25 (base 3/5) 9/25=

OpenStudy (rulnick):

The goal is to figure out what n would be for (3/5)^n to equal 9/25. So n=0 wouldn't work because (3/5)^0 = 1. And n=1 wouldn't either because (3/5)^1 = 3/5. So what power n would make (3/5)^n = 9/25?

OpenStudy (rulnick):

Hint: 3^2 = 9, and 5^2 = 25.

OpenStudy (anonymous):

@rulnick im lost

OpenStudy (rulnick):

How about (3/5)^2 (in other words, 3/5 raised to the second power)? The numerator would be 3^2 and the denominator would be 5^2. What would that make the resulting fraction?

OpenStudy (anonymous):

27/125 IS THT RIGHT

OpenStudy (rulnick):

(3/5)^3 would equal 27/125. But (3/5)^2 = 9/25. So the answer to your problem would be 9/25 = (3/5)^2.